# SIPDO: Closed-Loop Prompt Optimization via Synthetic Data Feedback

### v2

@@ Appendix A Optimzed Prompts for different tasks
In this section, we demonstrate optimized prompts by Chain-of-Thought (CoT), Automatic Prompt Engineering (APE), PromptAgent, and our method with Accuracys respectively.
@@
Table 4: Comparison of Optimized Prompts for Object Counting task, including CoT, APE, PromptAgent, and our method
@@
CoT Your task is to count the total number of objects mentioned in the question. Follow these simple steps to ensure accurate counting: **Steps to Follow:** 1. **Identify Items**: Read the question carefully and list all objects mentioned. 2. **Count Quantities**: For each item, check if a quantity is provided. If no quantity is mentioned, assume it is one. 3. **Add Totals**: Add up the quantities of all items to calculate the total count. 4. **Verify the Total**: Double-check to ensure no item is missed or counted twice. **Example:** - Question: "Count the apples, oranges, and bananas. There are 2 apples, 1 orange, and 3 bananas." - Step 1: Identify items: apples, oranges, bananas. - Step 2: Count quantities: 2 apples, 1 orange, 3 bananas. - Step 3: Add totals: 2 + 1 + 3 = 6. - Step 4: Verify: All items are accounted for, total is 6. - **Output**: "The total count is 6." Use this step-by-step method for every question to ensure accurate and clear results. 0.928
@@
Our Task Requirements: The task involves counting the total number of objects listed in a question. Each distinct object should be considered as part of the total count, regardless of its type or variation. The output should be formatted correctly as specified. Problem Rule Application: Identify all items listed in the question. Count each item exactly once, regardless of type, to determine the total number of objects. Ensure accuracy by verifying that all listed items have been included in the final count. Provide the final result in the required format: The number should be presented in both word form and numerical form, separated by a comma (e.g., "nine, 9"). No extra symbols, characters, or explanations should be included. Judgment Criteria: (Strictly follow these rules) Complete Identification: Extract and recognize all objects in the given list. Do not overlook any item mentioned in the question. Accurate Counting: Each item must be counted exactly once. Ensure no items are omitted or double-counted. Verification Process: Double-check the list to confirm that all objects are included. Cross-verify the final count to avoid errors. 0.923
Table 5: Comparison of Optimized Prompts for Penguins In A Table task, including CoT, APE, PromptAgent, and our method
@@
CoT You are tasked with answering questions about a table of penguins and their attributes. Use step-by-step reasoning to ensure accuracy in calculations and comparisons. The table is as follows: “‘ Name, Age, Height (cm), Weight (kg) Louis, 7, 50, 11 Bernard, 5, 80, 13 Vincent, 9, 60, 11 Gwen, 8, 70, 15 “‘ **Reasoning Steps for Each Question:** 1. Identify the target attribute (age, height, or weight) and the type of operation (comparison, ranking, filtering). 2. Extract the relevant rows or columns based on the question’s requirements. 3. Perform the required operation step-by-step using the extracted data. 4. Clearly summarize the answer based on the operation’s result. Example Workflow: - Question: "Who is the tallest penguin?" - Step 1: Identify the target attribute: Height. - Step 2: Extract the height values and corresponding names: [(Louis, 50), (Bernard, 80), (Vincent, 60), (Gwen, 70)]. - Step 3: Find the maximum height: Bernard (80 cm). - Step 4: Output the result: "Bernard is the tallest penguin with a height of 80 cm." Follow this workflow for every question to ensure clarity and correctness. 0.818
@@
PromptAgent Answer questions about a table of penguins and their attributes, considering both the penguin table and any additional relevant tables. Please provide step-by-step reasoning for your answers, and ensure to clarify any criteria used for filtering or sorting data. Here is a table where the first line is a header and each subsequent line is a penguin: name, age, height (cm), weight (kg) Louis, 7, 50, 11 Bernard, 5, 80, 13 Vincent, 9, 60, 11 Gwen, 8, 70, 15 For example: the age of Louis is 7, the weight of Gwen is 15 kg, the height of Bernard is 80 cm. What is the name of the last penguin sorted by alphabetic order? Options: (A) Louis (B) Bernard (C) Vincent (D) Gwen (E) James **Instructions**: 1. List the names of the penguins. 2. Sort the names alphabetically and present the sorted list clearly. 3. Identify the last name in the sorted list and indicate the corresponding option letter from the provided options. 4. If the last name does not match any of the options, select the name that is closest to the last name in the original list of penguins. At the end, show the answer option bracketed between <answer> and </answer>. 0.961
Our Answer questions about a dynamic, comprehensive table of penguins and their attributes that allows penguins and other animals to be added and removed. Perform calculations and comparisons based on the questions asked. Read the question carefully to determine which attribute is being compared (age, height, weight). When comparing an attribute, extract the name and that attribute, and then compare, ignoring the other attributes. Ensure the extracted value is from the correct column corresponding to the requested attribute. When using the table, align the data so that the first number is age, the second is height, and the third is weight. Understand the question correctly, find the key words from it, and then perform calculations or comparisons based on the key words The current table is as follows: Name, Age, Height (cm), Weight (kg) Louis, 7, 50, 11 Bernard, 5, 80, 13 Vincent, 9, 60, 11 Gwen, 8, 70, 15 **Question Rules to Apply:** - Identify the rows or columns that meet the specified conditions. - Retrieve the value of the required attribute from the identified rows or columns. When we modify this table (by adding new penguins or removing existing penguins or adding giraffes), we first confirm whether the information we added is a penguin or a giraffe, and then solve the problem of comparing, ranking, and filtering based on attributes between penguins or giraffes, depending on the problem. 0.964
Table 6: Comparison of Optimized Prompts for Geometric Shapes task, including CoT, APE, PromptAgent, and our method
@@
CoT Your task is to identify the geometric shape represented by the given SVG path data. Follow these steps to ensure accuracy: **Steps to Identify the Shape:** 1. **Check for ‘A’ Instructions**: – If the path contains ‘A’, determine: • **Circle**: 2 or more ‘A’ instructions. • **Sector**: 1 ‘A’ instruction. 2. **Count ‘L’ Instructions**: – If there are no ‘A’ instructions, count the ‘L’ instructions to determine the polygon’s shape: • **Line**: 2 ‘L’. • **Triangle**: 3 ‘L’. • **Rectangle**: 4 ‘L’. • **Pentagon**: 5 ‘L’. • **Hexagon**: 6 ‘L’. • **Heptagon**: 7 ‘L’. • **Octagon**: 8 ‘L’. • **Kite**: 4 ‘L’. 3. **Provide the Shape Name**: Output only the name of the shape (e.g., “circle”, “triangle”, “hexagon”). **Example:** – Input: ‘"M 10 10 L 20 10 L 20 20 L 10 20 Z"‘ – Step 1: No ‘A’ instructions. – Step 2: Count ‘L’ instructions: 4 ‘L’. – Step 3: Shape is a **Rectangle**. – **Output**: “rectangle”. Use this step-by-step process for all inputs to determine the correct shape. 0.791
@@
PromptAgent Analyze the SVG path data to identify the geometric shape it represents. Follow these comprehensive and refined steps to ensure accurate identification: 1. **Holistic Path Closure**: Determine if the path forms a closed shape by checking if the last point connects back to the starting point. If multiple ‘M‘ commands are present, analyze the segments collectively to identify any closed loops. Treat the entire path as a single entity for thorough analysis. 2. **Segment and Side Analysis**: Identify the types of segments used in the path: – **Line Segments**: Count the number of distinct line segments to determine the number of sides. Ensure accurate counting by considering all segments collectively. – **Arc Segments**: For paths using the ‘A‘ command, note that these represent elliptical arcs. Pay attention to the parameters to distinguish between circles and ellipses. 3. **In-depth Geometric Properties**: – For line segments, analyze the relative lengths of sides and angles between them. Consider properties such as parallel sides, equal side lengths, and right angles to distinguish between different types of polygons. Evaluate the overall shape formed by all segments. – For arc segments, examine the parameters of the ‘A‘ command: • Check if the radii are equal, which indicates a circle. • If the radii differ, consider the shape as an ellipse. 4. **Shape Identification and Classification**: Use the geometric properties to classify the shape: – For polygons, identify specific types like rectangles, kites, and trapezoids based on their properties. Pay special attention to the number of sides and the relationships between them. Consider the entire path as a single shape to ensure accurate classification. – For arcs, determine if the shape is a circle or an ellipse based on the radii. 5. **Options Selection and Interpretation**: Choose the most appropriate shape from the given options. Consider multiple interpretations of the path data, especially when multiple ‘M‘ commands are present, to ensure accurate classification. If the path represents multiple shapes, prioritize the most complex or relevant shape. 6. **Ambiguity Resolution**: In cases where the path data could represent multiple shapes, provide a rationale for selecting the most complex or relevant shape. Consider the context and any additional information that might influence the classification. 7. **Visual Verification**: If possible, visualize the path to confirm the identified shape. This step can help resolve any remaining ambiguities and ensure the accuracy of the classification. 8. **Iterative Refinement**: If the initial classification is uncertain, revisit the analysis steps to refine the identification. Consider alternative interpretations and re-evaluate the geometric properties. 9. **Contextual Considerations**: Take into account any contextual information or additional data that might influence the shape classification, especially in ambiguous cases. Provide your answer by selecting the correct option and enclosing it within ‘<answer>‘ and ‘</answer>‘ tags. **Example:** – SVG Path: ‘path d="M 8.10,55.86 L 1.74,25.57 L 12.08,23.40 L 18.44,53.69 L 8.10,55.86"‘ Analysis: The path forms a closed quadrilateral with opposite sides parallel and equal, indicating a rectangle. Answer: ‘<answer>H</answer>‘ – SVG Path: ‘path d="M 16.33,5.98 A 8.87,8.87 275.02 1,0 14.78,23.64 A 8.87,8.87 275.02 1,0 16.33,5.98"/‘ Analysis: The path uses elliptical arcs with equal radii, forming a closed loop, indicating a circle. Answer: ‘<answer>A</answer>‘ 0.830
@@
our Given the following SVG path data: “input” and options, identify the geometric shape it represents and provide **ONLY** the name of the shape as the ‘target’. **Task Requirements:** 1. Count the instructions in the SVG path 2. Judge the shape of the graphic according to the judgment criteria 3. Provide the exact name of the shape as output. You need to count how many instructions **L** are in the SVG path: **Problem Rule Application:** 1. Visualize the path data to understand the overall structure. 2. Find out whether there is instruction **A** in the instruction. If so, determine whether it is a circle or a sector according to the number of instructions **A**. If not, determine how many sides it is 3. For polygons, pay attention to the number of edges to identify the shape. The following are the number of instructions corresponding to different shapes: – **triangle**: 3 L – **rectangle**: 4 L – **hexagon**: 6 L – **pentagon**: 5 L – **octagon**: 8 L – **heptagon**: 7 L – **kite**: 4 L – **line**: 2 L – **circle**: 2 or more A – **sector**: 1 A **Judgment criteria:** (Please strictly abide by this rule) – No need to pay attention to “M” instructions – !! First identify whether there is an instruction “A” in the SVG path. If so, first determine whether it is a circle or a sector. – !! If there is no instruction “A”, determine the number of sides of the polygon based on the instruction “L”. A polygon with *n* sides requires *n* “L” instructions. (Please strictly abide by this rule) 0.822
@@
Table 7: Comparison of Optimized Prompts for Causal Judgment tasks, including CoT, APE, PromptAgent, and our method
@@
PromptAgent When addressing questions about causal attribution, ensure a comprehensive analysis by considering both individual and collective actions that contribute to an outcome. Clearly differentiate between necessary and sufficient causes, and recognize that multiple causes can simultaneously contribute to an outcome. Emphasize the importance of understanding both general and specific intentions, especially when outcomes are unintended. Define "intentional" actions as those where the actor or group had control over maintaining or altering the conditions necessary for the outcome, even if the specific result was not desired. Address scenarios where unintended consequences arise from intentional actions, and provide answers that reflect a nuanced understanding of how different elements interact to produce a result. Use diverse examples to illustrate key concepts like "direct causation," "simultaneity," and "unintended consequences," ensuring a balanced consideration of necessary and sufficient causes. Simplify complex scenarios by breaking them down into clear, manageable components, and provide definitions or examples of key terms to guide your analysis. Additionally, clarify definitions of key terms such as "necessary," "sufficient," "intentional," and "unintended consequences" to ensure precise understanding. Highlight the importance of interactions between multiple causes, especially in complex scenarios, and offer strategies for analyzing scenarios where simultaneity is crucial. Explore the nuances of intentional actions and unintended consequences more deeply, and encourage the use of diverse examples to illustrate different aspects of causation. Pay special attention to the role of individual actions in maintaining necessary conditions and the distinction between collective and individual causation. Emphasize that in collective decision-making, the outcome can be intentional if it aligns with the group’s goals, even if individual members disagree. Reinforce the distinction between necessary and sufficient causes, ensuring the model understands that necessary causes alone do not determine the outcome. Clarify that following a protocol does not remove intentionality if the outcome aligns with organizational priorities. Highlight that intentionality can be attributed if the outcome was a foreseeable consequence of the action, regardless of individual opposition. 0.846
@@
Our Task Requirements Determine whether a given event (cause) directly leads to another event (effect). Assess the causal relationship based on logical reasoning, ensuring a clear and definitive answer. The final output must be only "Yes" or "No", strictly adhering to the required format. Problem Rule Application Identify the cause and effect within the question. Assess necessity: Determine if the cause is essential for the effect to occur. Evaluate causation: If the cause did not happen, would the effect still occur? If the effect only happens when the cause is present, then the cause directly leads to the effect. If the effect can still happen independently, then the relationship is not causal. Judgment Criteria Direct Causation: If the cause directly leads to the effect and is a necessary condition, answer "Yes". If the effect would not have occurred without the cause, answer "Yes". Example: "Dropping a glass caused it to shatter." → Yes. No Direct Causation: If the effect can occur without the cause, answer "No". If the cause is only correlated but not necessary, answer "No". Example: "Wearing a red shirt caused the stock market to rise." → No. Verification Process: Check whether the absence of the cause results in the absence of the effect. Ensure logical consistency in the causal assessment. 0.880
@@
Table 8: Comparison of Optimized Prompts for Epistemic task, including CoT, APE, PromptAgent, and our method
@@
PromptAgent Determine the relationship between two sentences by evaluating whether the first sentence provides direct or logically implied evidence for the second. Choose from the options ’entailment’ or ’non-entailment’. Consider the following: - **Entailment**: The first sentence directly or through logical implication confirms the truth of the second sentence, even if it involves a chain of beliefs or perceptions, as long as the chain logically supports the hypothesis. - **Non-entailment**: The first sentence does not confirm the truth of the second sentence, often involving unsupported assumptions, beliefs, or suspicions that do not logically lead to the hypothesis. Guidelines for Analysis: 1. **Clarify Belief Chains and Logical Implications**: Understand how belief chains work and when they logically support the hypothesis. Pay attention to verbs indicating beliefs, assumptions, or suspicions (e.g., "thinks," "assumes," "suspects") versus those indicating direct evidence (e.g., "learns," "knows," "remembers"). Consider how these verbs interact in belief chains and what they imply about the subject’s own beliefs. 2. **Evaluate Direct and Implied Evidence**: Determine if the premise provides direct or logically implied evidence for the hypothesis, considering how belief chains can logically support the hypothesis. Recognize that indirect beliefs about another person’s recognition can imply one’s own belief about a situation, especially when the belief chain is logical and straightforward. 3. **Consider Perspective and Source of Information**: Note any differences in perspective or source of information (e.g., who remembers or assumes something) and how these perspectives contribute to the logical implication of the hypothesis. 4. **Conduct a Comprehensive Analysis**: Use a step-by-step approach to ensure all relevant details and logical implications are considered in the analysis. Balance the emphasis on direct evidence with the recognition of logical implications from indirect beliefs. Example: Premise: "Charlotte thinks that Richard recognizes that a boy is standing in a pool getting splashed with water." Hypothesis: "Charlotte thinks that a boy is standing in a pool getting splashed with water." Options: (A) entailment (B) non-entailment Analysis: 1. **Understanding the Premise**: The premise indicates that Charlotte thinks Richard recognizes a specific situation involving a boy in a pool. 2. **Understanding the Hypothesis**: The hypothesis states that Charlotte thinks a boy is in a pool getting splashed with water. 3. **Assessing the Relationship**: The premise implies that Charlotte has a belief about the situation (through Richard’s recognition), which logically supports the hypothesis. Charlotte’s belief about Richard’s recognition suggests she also believes in the situation’s occurrence. 4. **Conclusion**: The relationship is one of entailment because Charlotte’s belief about Richard’s recognition logically implies her belief in the situation. Therefore, the correct answer is: <answer>A</answer> Identify the relation between the following premises and hypotheses, choosing from the options ’entailment’ or ’non-entailment’. At the end, show the answer option bracketed between <answer> and </answer>. 0.916

### v3

@@ Appendix B Optimzed Prompts for different tasks
In this section, we demonstrate optimized prompts by Chain-of-Thought (CoT), Automatic Prompt Engineering (APE), PromptAgent, and SIPDO with Accuracys respectively.
@@
Table 9: Comparison of Optimized Prompts for Object Counting task, including CoT, APE, PromptAgent, and SIPDO
@@
CoT Your task is to count the total number of objects mentioned in the question. Follow these simple steps to ensure accurate counting: **Steps to Follow:** 1. **Identify Items**: Read the question carefully and list all objects mentioned. 2. **Count Quantities**: For each item, check if a quantity is provided. If no quantity is mentioned, assume it is one. 3. **Add Totals**: Add up the quantities of all items to calculate the total count. 4. **Verify the Total**: Double-check to ensure no item is missed or counted twice. **Example:** - Question: ”Count the apples, oranges, and bananas. There are 2 apples, 1 orange, and 3 bananas.” - Step 1: Identify items: apples, oranges, bananas. - Step 2: Count quantities: 2 apples, 1 orange, 3 bananas. - Step 3: Add totals: 2 + 1 + 3 = 6. - Step 4: Verify: All items are accounted for, total is 6. - **Output**: ”The total count is 6.” Use this step-by-step method for every question to ensure accurate and clear results. 0.928
@@
SPIDO Task Requirements: The task involves counting the total number of objects listed in a question. Each distinct object should be considered as part of the total count, regardless of its type or variation. The output should be formatted correctly as specified. Problem Rule Application: Identify all items listed in the question. Count each item exactly once, regardless of type, to determine the total number of objects. Ensure accuracy by verifying that all listed items have been included in the final count. Provide the final result in the required format: The number should be presented in both word form and numerical form, separated by a comma (e.g., ”nine, 9”). No extra symbols, characters, or explanations should be included. Judgment Criteria: (Strictly follow these rules) Complete Identification: Extract and recognize all objects in the given list. Do not overlook any item mentioned in the question. Accurate Counting: Each item must be counted exactly once. Ensure no items are omitted or double-counted. Verification Process: Double-check the list to confirm that all objects are included. Cross-verify the final count to avoid errors. 0.923
Table 10: Comparison of Optimized Prompts for Penguins In A Table task, including CoT, APE, PromptAgent, and SIPDO
@@
CoT You are tasked with answering questions about a table of penguins and their attributes. Use step-by-step reasoning to ensure accuracy in calculations and comparisons. The table is as follows: “‘ Name, Age, Height (cm), Weight (kg) Louis, 7, 50, 11 Bernard, 5, 80, 13 Vincent, 9, 60, 11 Gwen, 8, 70, 15 “‘ **Reasoning Steps for Each Question:** 1. Identify the target attribute (age, height, or weight) and the type of operation (comparison, ranking, filtering). 2. Extract the relevant rows or columns based on the question’s requirements. 3. Perform the required operation step-by-step using the extracted data. 4. Clearly summarize the answer based on the operation’s result. Example Workflow: - Question: ”Who is the tallest penguin?” - Step 1: Identify the target attribute: Height. - Step 2: Extract the height values and corresponding names: [(Louis, 50), (Bernard, 80), (Vincent, 60), (Gwen, 70)]. - Step 3: Find the maximum height: Bernard (80 cm). - Step 4: Output the result: ”Bernard is the tallest penguin with a height of 80 cm.” Follow this workflow for every question to ensure clarity and correctness. 0.818
@@
PromptAgent Answer questions about a table of penguins and their attributes, considering both the penguin table and any additional relevant tables. Please provide step-by-step reasoning for your answers, and ensure to clarify any criteria used for filtering or sorting data. Here is a table where the first line is a header and each subsequent line is a penguin: name, age, height (cm), weight (kg) Louis, 7, 50, 11 Bernard, 5, 80, 13 Vincent, 9, 60, 11 Gwen, 8, 70, 15 For example: the age of Louis is 7, the weight of Gwen is 15 kg, the height of Bernard is 80 cm. What is the name of the last penguin sorted by alphabetic order? Options: (A) Louis (B) Bernard (C) Vincent (D) Gwen (E) James **Instructions**: 1. List the names of the penguins. 2. Sort the names alphabetically and present the sorted list clearly. 3. Identify the last name in the sorted list and indicate the corresponding option letter from the provided options. 4. If the last name does not match any of the options, select the name that is closest to the last name in the original list of penguins. At the end, show the answer option bracketed between ¡answer¿ and ¡/answer¿. 0.961
SIPDO Answer questions about a dynamic, comprehensive table of penguins and their attributes that allows penguins and other animals to be added and removed. Perform calculations and comparisons based on the questions asked. Read the question carefully to determine which attribute is being compared (age, height, weight). When comparing an attribute, extract the name and that attribute, and then compare, ignoring the other attributes. Ensure the extracted value is from the correct column corresponding to the requested attribute. When using the table, align the data so that the first number is age, the second is height, and the third is weight. Understand the question correctly, find the key words from it, and then perform calculations or comparisons based on the key words The current table is as follows: Name, Age, Height (cm), Weight (kg) Louis, 7, 50, 11 Bernard, 5, 80, 13 Vincent, 9, 60, 11 Gwen, 8, 70, 15 **Question Rules to Apply:** - Identify the rows or columns that meet the specified conditions. - Retrieve the value of the required attribute from the identified rows or columns. When we modify this table (by adding new penguins or removing existing penguins or adding giraffes), we first confirm whether the information we added is a penguin or a giraffe, and then solve the problem of comparing, ranking, and filtering based on attributes between penguins or giraffes, depending on the problem. 0.964
Table 11: Comparison of Optimized Prompts for Geometric Shapes task, including CoT, APE, PromptAgent, and SIPDO
@@
CoT Your task is to identify the geometric shape represented by the given SVG path data. Follow these steps to ensure accuracy: **Steps to Identify the Shape:** 1. **Check for ‘A’ Instructions**: – If the path contains ‘A’, determine: • **Circle**: 2 or more ‘A’ instructions. • **Sector**: 1 ‘A’ instruction. 2. **Count ‘L’ Instructions**: – If there are no ‘A’ instructions, count the ‘L’ instructions to determine the polygon’s shape: • **Line**: 2 ‘L’. • **Triangle**: 3 ‘L’. • **Rectangle**: 4 ‘L’. • **Pentagon**: 5 ‘L’. • **Hexagon**: 6 ‘L’. • **Heptagon**: 7 ‘L’. • **Octagon**: 8 ‘L’. • **Kite**: 4 ‘L’. 3. **Provide the Shape Name**: Output only the name of the shape (e.g., “circle”, “triangle”, “hexagon”). **Example:** – Input: ‘”M 10 10 L 20 10 L 20 20 L 10 20 Z”‘ – Step 1: No ‘A’ instructions. – Step 2: Count ‘L’ instructions: 4 ‘L’. – Step 3: Shape is a **Rectangle**. – **Output**: “rectangle”. Use this step-by-step process for all inputs to determine the correct shape. 0.791
@@
PromptAgent Analyze the SVG path data to identify the geometric shape it represents. Follow these comprehensive and refined steps to ensure accurate identification: 1. **Holistic Path Closure**: Determine if the path forms a closed shape by checking if the last point connects back to the starting point. If multiple ‘M‘ commands are present, analyze the segments collectively to identify any closed loops. Treat the entire path as a single entity for thorough analysis. 2. **Segment and Side Analysis**: Identify the types of segments used in the path: – **Line Segments**: Count the number of distinct line segments to determine the number of sides. Ensure accurate counting by considering all segments collectively. – **Arc Segments**: For paths using the ‘A‘ command, note that these represent elliptical arcs. Pay attention to the parameters to distinguish between circles and ellipses. 3. **In-depth Geometric Properties**: – For line segments, analyze the relative lengths of sides and angles between them. Consider properties such as parallel sides, equal side lengths, and right angles to distinguish between different types of polygons. Evaluate the overall shape formed by all segments. – For arc segments, examine the parameters of the ‘A‘ command: • Check if the radii are equal, which indicates a circle. • If the radii differ, consider the shape as an ellipse. 4. **Shape Identification and Classification**: Use the geometric properties to classify the shape: – For polygons, identify specific types like rectangles, kites, and trapezoids based on their properties. Pay special attention to the number of sides and the relationships between them. Consider the entire path as a single shape to ensure accurate classification. – For arcs, determine if the shape is a circle or an ellipse based on the radii. 5. **Options Selection and Interpretation**: Choose the most appropriate shape from the given options. Consider multiple interpretations of the path data, especially when multiple ‘M‘ commands are present, to ensure accurate classification. If the path represents multiple shapes, prioritize the most complex or relevant shape. 6. **Ambiguity Resolution**: In cases where the path data could represent multiple shapes, provide a rationale for selecting the most complex or relevant shape. Consider the context and any additional information that might influence the classification. 7. **Visual Verification**: If possible, visualize the path to confirm the identified shape. This step can help resolve any remaining ambiguities and ensure the accuracy of the classification. 8. **Iterative Refinement**: If the initial classification is uncertain, revisit the analysis steps to refine the identification. Consider alternative interpretations and re-evaluate the geometric properties. 9. **Contextual Considerations**: Take into account any contextual information or additional data that might influence the shape classification, especially in ambiguous cases. Provide your answer by selecting the correct option and enclosing it within ‘¡answer¿‘ and ‘¡/answer¿‘ tags. **Example:** – SVG Path: ‘path d=”M 8.10,55.86 L 1.74,25.57 L 12.08,23.40 L 18.44,53.69 L 8.10,55.86”‘ Analysis: The path forms a closed quadrilateral with opposite sides parallel and equal, indicating a rectangle. Answer: ‘¡answer¿H¡/answer¿‘ – SVG Path: ‘path d=”M 16.33,5.98 A 8.87,8.87 275.02 1,0 14.78,23.64 A 8.87,8.87 275.02 1,0 16.33,5.98”/‘ Analysis: The path uses elliptical arcs with equal radii, forming a closed loop, indicating a circle. Answer: ‘¡answer¿A¡/answer¿‘ 0.830
@@
SIPDO Given the following SVG path data: “input” and options, identify the geometric shape it represents and provide **ONLY** the name of the shape as the ‘target’. **Task Requirements:** 1. Count the instructions in the SVG path 2. Judge the shape of the graphic according to the judgment criteria 3. Provide the exact name of the shape as output. You need to count how many instructions **L** are in the SVG path: **Problem Rule Application:** 1. Visualize the path data to understand the overall structure. 2. Find out whether there is instruction **A** in the instruction. If so, determine whether it is a circle or a sector according to the number of instructions **A**. If not, determine how many sides it is 3. For polygons, pay attention to the number of edges to identify the shape. The following are the number of instructions corresponding to different shapes: – **triangle**: 3 L – **rectangle**: 4 L – **hexagon**: 6 L – **pentagon**: 5 L – **octagon**: 8 L – **heptagon**: 7 L – **kite**: 4 L – **line**: 2 L – **circle**: 2 or more A – **sector**: 1 A **Judgment criteria:** (Please strictly abide by this rule) – No need to pay attention to “M” instructions – !! First identify whether there is an instruction “A” in the SVG path. If so, first determine whether it is a circle or a sector. – !! If there is no instruction “A”, determine the number of sides of the polygon based on the instruction “L”. A polygon with *n* sides requires *n* “L” instructions. (Please strictly abide by this rule) 0.822
@@
Table 12: Comparison of Optimized Prompts for Causal Judgment tasks, including CoT, APE, PromptAgent, and SIPDO
@@
PromptAgent When addressing questions about causal attribution, ensure a comprehensive analysis by considering both individual and collective actions that contribute to an outcome. Clearly differentiate between necessary and sufficient causes, and recognize that multiple causes can simultaneously contribute to an outcome. Emphasize the importance of understanding both general and specific intentions, especially when outcomes are unintended. Define ”intentional” actions as those where the actor or group had control over maintaining or altering the conditions necessary for the outcome, even if the specific result was not desired. Address scenarios where unintended consequences arise from intentional actions, and provide answers that reflect a nuanced understanding of how different elements interact to produce a result. Use diverse examples to illustrate key concepts like ”direct causation,” ”simultaneity,” and ”unintended consequences,” ensuring a balanced consideration of necessary and sufficient causes. Simplify complex scenarios by breaking them down into clear, manageable components, and provide definitions or examples of key terms to guide your analysis. Additionally, clarify definitions of key terms such as ”necessary,” ”sufficient,” ”intentional,” and ”unintended consequences” to ensure precise understanding. Highlight the importance of interactions between multiple causes, especially in complex scenarios, and offer strategies for analyzing scenarios where simultaneity is crucial. Explore the nuances of intentional actions and unintended consequences more deeply, and encourage the use of diverse examples to illustrate different aspects of causation. Pay special attention to the role of individual actions in maintaining necessary conditions and the distinction between collective and individual causation. Emphasize that in collective decision-making, the outcome can be intentional if it aligns with the group’s goals, even if individual members disagree. Reinforce the distinction between necessary and sufficient causes, ensuring the model understands that necessary causes alone do not determine the outcome. Clarify that following a protocol does not remove intentionality if the outcome aligns with organizational priorities. Highlight that intentionality can be attributed if the outcome was a foreseeable consequence of the action, regardless of individual opposition. 0.846
@@
SIPDO Task Requirements Determine whether a given event (cause) directly leads to another event (effect). Assess the causal relationship based on logical reasoning, ensuring a clear and definitive answer. The final output must be only ”Yes” or ”No”, strictly adhering to the required format. Problem Rule Application Identify the cause and effect within the question. Assess necessity: Determine if the cause is essential for the effect to occur. Evaluate causation: If the cause did not happen, would the effect still occur? If the effect only happens when the cause is present, then the cause directly leads to the effect. If the effect can still happen independently, then the relationship is not causal. Judgment Criteria Direct Causation: If the cause directly leads to the effect and is a necessary condition, answer ”Yes”. If the effect would not have occurred without the cause, answer ”Yes”. Example: ”Dropping a glass caused it to shatter.” → Yes. No Direct Causation: If the effect can occur without the cause, answer ”No”. If the cause is only correlated but not necessary, answer ”No”. Example: ”Wearing a red shirt caused the stock market to rise.” → No. Verification Process: Check whether the absence of the cause results in the absence of the effect. Ensure logical consistency in the causal assessment. 0.880
@@
Table 13: Comparison of Optimized Prompts for Epistemic task, including CoT, APE, PromptAgent, and SIPDO
@@
PromptAgent Determine the relationship between two sentences by evaluating whether the first sentence provides direct or logically implied evidence for the second. Choose from the options ’entailment’ or ’non-entailment’. Consider the following: - **Entailment**: The first sentence directly or through logical implication confirms the truth of the second sentence, even if it involves a chain of beliefs or perceptions, as long as the chain logically supports the hypothesis. - **Non-entailment**: The first sentence does not confirm the truth of the second sentence, often involving unsupported assumptions, beliefs, or suspicions that do not logically lead to the hypothesis. Guidelines for Analysis: 1. **Clarify Belief Chains and Logical Implications**: Understand how belief chains work and when they logically support the hypothesis. Pay attention to verbs indicating beliefs, assumptions, or suspicions (e.g., ”thinks,” ”assumes,” ”suspects”) versus those indicating direct evidence (e.g., ”learns,” ”knows,” ”remembers”). Consider how these verbs interact in belief chains and what they imply about the subject’s own beliefs. 2. **Evaluate Direct and Implied Evidence**: Determine if the premise provides direct or logically implied evidence for the hypothesis, considering how belief chains can logically support the hypothesis. Recognize that indirect beliefs about another person’s recognition can imply one’s own belief about a situation, especially when the belief chain is logical and straightforward. 3. **Consider Perspective and Source of Information**: Note any differences in perspective or source of information (e.g., who remembers or assumes something) and how these perspectives contribute to the logical implication of the hypothesis. 4. **Conduct a Comprehensive Analysis**: Use a step-by-step approach to ensure all relevant details and logical implications are considered in the analysis. Balance the emphasis on direct evidence with the recognition of logical implications from indirect beliefs. Example: Premise: ”Charlotte thinks that Richard recognizes that a boy is standing in a pool getting splashed with water.” Hypothesis: ”Charlotte thinks that a boy is standing in a pool getting splashed with water.” Options: (A) entailment (B) non-entailment Analysis: 1. **Understanding the Premise**: The premise indicates that Charlotte thinks Richard recognizes a specific situation involving a boy in a pool. 2. **Understanding the Hypothesis**: The hypothesis states that Charlotte thinks a boy is in a pool getting splashed with water. 3. **Assessing the Relationship**: The premise implies that Charlotte has a belief about the situation (through Richard’s recognition), which logically supports the hypothesis. Charlotte’s belief about Richard’s recognition suggests she also believes in the situation’s occurrence. 4. **Conclusion**: The relationship is one of entailment because Charlotte’s belief about Richard’s recognition logically implies her belief in the situation. Therefore, the correct answer is: <answer>A</answer> Identify the relation between the following premises and hypotheses, choosing from the options ’entailment’ or ’non-entailment’. At the end, show the answer option bracketed between <answer> and </answer>. 0.916